1. Show that if 5/3< 2x< 11/3, then x∈{y∈R such that |y- 4/3| < 1/2}


2. Draw the graph of the function f, defined by f(x)= |x-6| + |5 - x| ; x∈ [2,8]


1
Expert's answer
2022-06-03T06:21:12-0400

1.

Let's solve the inequalities.

"\frac{5}{3}< 2x< \frac{11}{3},"

we can divide by 2:

"\frac{5}{6}<x<\frac{11}{6},"

in interval notation: "x \in (\frac{5}{6}, \frac{11}{6})."


"|y- \frac{4}{3}| < \frac{1}{2},"

"- \frac{1}{2}<y- \frac{4}{3} < \frac{1}{2},"

"- \frac{1}{2}+ \frac{4}{3}<y < \frac{1}{2}+ \frac{4}{3},"

"- \frac{3}{6}+ \frac{8}{6}<y < \frac{3}{6}+ \frac{8}{6},"

"\frac{5}{6}<y<\frac{11}{6},"

in interval notation: "y \in (\frac{5}{6}, \frac{11}{6})."


We obtain the same intervals, so the first statement holds.

2.

"f(x)= |x-6| + |5 - x| ; x\u2208 [2,8]"

Let's consider this function on three intervals:

"f(x)=-(x-6) + (5 - x)=-x+6+5-x="

"=-2x+11; x\u2208 [2,5)"


"f(x)=-(x-6) - (5 - x)=-x+6-5+x=1, x\in[5, 6)"


"f(x)=(x-6) - (5 - x)=x-6-5+x="

"=2x-11, x\in[6, 8]"


Now we can draw the graph of the function f: